4x+5y+3z=15
X-3y+2z=-6
-x+2y-z=3

Answer by TimothyLamb(4379)

---
4x+5y+3z=15
x-3y+2z=-6
-x+2y-z=3
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 2
y = 2
z = -1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php